OpenStudy (anonymous):

Solve for z: -2+(z-5)^2/3=6

OpenStudy (anonymous):

1) add 2 2) cube both sides 3) take the square root

OpenStudy (anonymous):

is this \[-2+(z-5)^{\frac{2}{3}}=6\] or \[-2+(z-5)^{\frac{3}{2}}=6\]

OpenStudy (anonymous):

1) add 2 2) multiply by 3 3) square root (remember + and -) 4) add 5

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

@satellite73 powers have a higher order than division

OpenStudy (anonymous):

aaaaand the asker mistyping the question strikes again!

OpenStudy (anonymous):

mistyping the question? how so....

OpenStudy (anonymous):

If you wanted it the same as satellite's first equation you needed a bracket around the 2/3

OpenStudy (anonymous):

There is no brackets in the original problem so therefore I can not put it on there if it is not there. That then would be mistyping the question.

OpenStudy (anonymous):

you didnt use superscripts though which would be present in your question. using brackets stops the need for superscripts.

OpenStudy (anonymous):

\[-2+(z-5)^{\frac{2}{3}}=6\] \[(z-5)^{\frac{2}{3}}=8\] \[(z-5)^2=8^3=512\] \[z-5=\pm\sqrt{512}=\pm16\sqrt{2}\] \[z=5\pm16\sqrt{2}\]

OpenStudy (anonymous):

that is if it was \[-2+(z-5)^{\frac{2}{3}}=6\]

OpenStudy (anonymous):

if it was \[-2+(z-5)^{\frac{3}{2}}=6\] then \[(z-5)^{\frac{3}{2}}=8\] \[\sqrt{z-5}=2\] \[z-5=4\] \[z=9\]

OpenStudy (anonymous):

Satellite; thank you for your help. I got the same answer I was working on it. :)